### v2

# A Fourth-Order Cut-cell Multigrid Method

Abstract
To numerically solve a generic elliptic equation on two-dimensional domains with rectangular Cartesian grids, we propose a cut-cell geometric multigrid method that features (1) general algorithmic steps that apply to all forms of elliptic equations and all types of boundary conditions, (2) the versatility of handling both regular and irregular domains with arbitrarily complex topology and geometry, (3) the fourth-order accuracy even at the presence of ${\cal C}^{1}$ discontinuities on the domain boundary, and (4) the optimal complexity of $O(h^{-2})$. Test results demonstrate the generality, accuracy, efficiency, robustness, and excellent conditioning of the proposed method.
keywords:
elliptic equations, geometric multigrid methods, finite-volume methods, poised lattice generation, generating cut cells without small volumes.
{AMS}
65N08, 65N55
1 Introduction
Consider the constant-coefficient elliptic equation
(1a) ${\cal L}u$ $:=a\frac{\partial^{2}u}{\partial x^{2}}+b\frac{\partial^{2}u}{\partial x\partial y}+c\frac{\partial^{2}u}{\partial y^{2}}=f(x,y)\quad$ $\text{in }\Omega,$
(1b) $\mathcal{N}u$ $=g(x,y)$ $\text{on }\partial\Omega.$
where the domain $\Omega$ is an open subset of $\mathbb{R}^{2}$, $u$ the unknown function, $a,b,c$ real numbers satisfying $b^{2}-4ac<0$, and $\mathcal{N}$ the boundary-condition operator:
[…]
with $\alpha_{1},\alpha_{2}\in\mathbb{R}$ and $\mathbf{I}_{\text{d}}$ as the identity operator.
Set $(a,b,c)=(1,0,1)$ in (1) and we get Poisson’s equation, which is particularly important for designing numerical methods for solving partial differential equations (PDEs). For example, at the core of the projection methods [5, 18, 26, 37, 38, 21] for the incompressible Navier–Stokes equations (INSE) is the numerical solution of a sequence of Poisson’s equations and Helmholtz-like equations.
A myriad of numerical methods have been developed for solving PDEs on regular domains. In many real-world applications, however, a problem domain may have complex topology and irregular geometry. Within the realm of finite element methods (FEMs), the irregular geometry is handled either by an external mesh generator, yielding the interface-fitted FEMs [2, 3, 6, 30], or by local treatments of the irregular boundary inside its algorithms, leading to the interface-unfitted FEMs [23, 13, 14]. As for finite difference (FD) methods, two notable examples are the immersed interface method [20, 22, 24, 25], where the discretization stencil is modified to incorporate jump conditions on the interface, and the ghost fluid method [11, 27, 28, 35], where physical variables are smoothly extended across the interface so that conventional FD formulas can be used for spatial discretization. For other FD methods on irregular domains, see [42, 15, 8, 31, 41] and references therein.
For finite volume (FV) formulations, the cut cell method, also known as the embedded boundary (EB) method, consists of three main steps as follows.
  1. (CCM-1)
    Use a Cartesian grid to partition the domain $\Omega$ into a set of cut cells, which are irregular near an irregular boundary and regular otherwise.
  2. (CCM-2)
    Approximate the average of ${\cal L}u$ over each cut cell by a linear expression of averages of $u$ over nearby cut cells, cf. (10), and consequently discretize (1) into a system of linear equations.
  3. (CCM-3)
    Solve the linear system to obtain averages of $u$ over all cut cells as the numerical solution of (1).
For Poisson’s equation in two-dimensions (2D), Johansen and Colella [17] proposed a second-order EB method, in which the average of $\nabla\cdot(\beta\nabla u)$ over a cut cell is transformed by the divergence theorem into a sum of fluxes through cell faces. The fluxes through regular faces are approximated by standard FV formulas while those near an irregular domain boundary by quadratic interpolations. This second-order EB method has been extended to the three-dimensional Poisson’s equation [32], the heat equation [29, 32], and the INSE [19, 33]. To improve the second-order accuracy to the fourth order, one must successfully address three main difficulties, viz. the representation of irregular geometry, the approximation of integrals over irregular cut cells, and the discretization of (1) with sufficient accuracy.
As far as fourth-order FV methods on irregular domains are concerned, we are only aware of the high-order EB method developed by Collela [7] and colleagues [9], in which the irregular domain is represented as $\Omega=\{\mathbf{x}:\phi(\mathbf{x})<0\}$ with $\phi$ being a smooth level set function $\mathbb{R}^{2}\rightarrow\mathbb{R}$, the domain boundary $\partial\Omega$ as $\phi^{-1}(0)$, and the unit normal vector of $\partial\Omega$ as $\mathbf{n}=\frac{\nabla\phi}{\|\nabla\phi\|}$. Assuming the existence of a flux vector $\mathbf{F}(u)$ satisfying ${\cal L}u=\nabla\cdot\mathbf{F}(u)$, they use the divergence theorem to transform the integral of ${\cal L}u$ over an irregular cut cell to those of $\mathbf{F}(u)$ over the cell faces. To further discretize (1) on a cut cell, they expand $\mathbf{F}(u)$ in its Taylor series, fit a local multivariate polynomial via weighted least squares [9, §2.3], calculate moments of monomials over the cell faces, and finally approximate the integral of $\nabla\cdot\mathbf{F}(u)$ over the cut cell with a linear combination of integrals of $\mathbf{F}(u)$ over nearby cell faces. As such, the approximation of the irregular geometry and the discretization of (1) are tightly coupled.
In spite of its successes, the aforementioned fourth-order EB method [7, 9] has a number of limitations. First, the assumption of the divergence form ${\cal L}u=\nabla\cdot\mathbf{F}(u)$ limits the generality of the EB method, since many elliptic equations with Neumann conditions and a cross-derivative term have no divergence form. Second, the representation of irregular geometry by a level set function $\phi$ leads to severe accuracy deterioration at the presence of kinks, i.e., ${\cal C}^{1}$ discontinuities, at which $\frac{\nabla\phi}{\|\nabla\phi\|}$ has an $O(1)$ error in approximating the normal vector $\mathbf{n}$. Indeed, Devendran et al. [9] reported that the accuracy of their fourth-order EB method drops to the first order or non-convergence in solving a Poisson equation on a domain with kinks, cf. Table 4. Although this accuracy deterioration can be alleviated by mollifying the kinks, this mollification is not applicable to all cases, and its effectiveness depends heavily on the nature of the equation and the mollification formula [9]. Third, there is no guarantee that the lattices (or stencils) for weighted least squares such as that in [9, Sec. 2.3.3] work well for all geometry. On one hand, a wide stencil may lead to a large number of redundant points, adversely affecting computational efficiency. Then it is desirable to have a poised lattice whose cardinality equals the dimension of the space of multivariate interpolating polynomials [41]. On the other hand, the local geometry might make it impossible to fit a high-order multivariate polynomial out of nearby cut cells. In this case, one needs to know the highest degree of interpolating polynomials for which the local geometry admits. As far as we know, the only algorithm that meets these requirements is that of the poised lattice generation (PLG) [41] for FD methods.
The above discussions pertain to (CCM-1) and (CCM-2). As for (CCM-3), the linear system that results from discretizing the elliptic equation is typically solved by a geometric multigrid method on regular domains and by an algebraic multigrid method on irregular domains [4]. Some researchers [9, 17, 32] extend geometric multigrid methods to irregular boundaries by modifying the restriction and interpolation operators on cells near irregular boundaries; but it is not clear whether these multigrid methods can achieve the optimal complexity of $O(h^{-2})$. It is observed in [33] that these methods struggle with convergence on domains with very complex geometry.
The above discussions lead to questions as follows.
  1. (Q-1)
    To represent an irregular domain $\Omega$ with arbitrary geometry and topology, can we have a simple and efficient scheme that is always fourth-order accurate? Furthermore, for a given threshold $\epsilon\in(0,1)$ and a Cartesian grid of size $h$, can we partition $\Omega$ into a set of cut cells whose volumes are between $\epsilon h^{2}$ and $2h^{2}$?
  2. (Q-2)
    Can we design an FV discretization of (1) on these cut cells so that (i) the discretization processes depends neither on values of $(a,b,c)$ in (1a) nor on forms of boundary conditions in (1b), and (ii) the fourth-order accuracy depends neither on the topology of $\Omega$ nor on the absence of kinks on $\partial\Omega$?
  3. (Q-3)
    For the FV discretization in (Q-2), can we further develop a geometric multigrid method that solves the resulting linear system with optimal complexity?
In this paper, we give positive answers to all above questions by proposing a cut-cell geometric multigrid method for solving (1) over arbitrary 2D domains.
Our answer to (Q-1) is based on Yin sets [40], a mathematical model of 2D continua with arbitrary topology, which we briefly review in Section 2. Utilizing the Boolean algebra of Yin sets in [40], we propose a cut-cell algorithm in Section 3 to generate a set $\mathsf{C}_{\epsilon}^{h}$ of cut cells so that the regularized union of these cut cells equals $\Omega$ and the volume of each cut cell is no less than $\epsilon h^{2}$, precluding the well-known small-volume problem in FV methods.
(Q-2) is answered in Section 4. We call a cut cell a symmetric finite volume (SFV) cell if classical symmetric FV formulas apply to it; otherwise it is called a PLG cell, cf. Definition 4.1. The fourth-order discretization of the integral of (1) over SFV cells is given in Section 4.1 and that over PLG cells is based on the PLG algorithm [41] summarized in Section 4.2.1. Given $K\subset\mathbb{Z}^{\scriptsize\textsf{D}}$, a starting point $q\in K$, and the total degree $n$ of D-variate polynomials, this PLG algorithm generates a poised lattice on which D-variate polynomial interpolation is unisolvent. In Section 4.2.2, this PLG algorithm is adapted to the FV formulation to generate linear equations that approximates integrals of (1) over PLG cells to sufficient accuracy. The complete linear system is summarized in Section 4.3.
In Section 5, we answer (Q-3) by proposing a cut-cell geometric multigrid method, which hinges on the fact that numbers of PLG and SFV cells are $O(h^{-1})$ and $O(h^{-2})$, respectively; see the opening paragraph of Section 5 for other key ideas. In Section 6, we demonstrate the accuracy, efficiency, generality, robustness, and excellent conditioning of the proposed cut-cell method by results of a number of numerical tests. We conclude this work in Section 7 with several future research prospects.
2 Modeling continua with Yin sets
In this section, we briefly review Yin sets [40] as a model of topological structures and geometric features of 2D continua.
In a topological space $\mathcal{X}$}, the complement of a subset ${\mathcal{P}}\subseteq{\mathcal{X}}$, written $\mathcal{P}$$^{\prime}$, is the set ${\mathcal{X}}\setminus{\mathcal{P}}$. The closure of a set ${\mathcal{P}}\subseteq{\mathcal{X}}$, written $\overline{\mathcal{P}}$, is the intersection of all closed supersets of $\mathcal{P}$}. The interior of $\mathcal{P}$}, written $\mathcal{P}$$^{\circ}$, is the union of all open subsets of $\mathcal{P}$}. The exterior of $\mathcal{P}$}, written ${\mathcal{P}}^{\perp}:={\mathcal{P}}^{\prime\circ}:=({\mathcal{P}}^{\prime})^{\circ}$, is the interior of its complement. A point $\mathbf{x}\in{\mathcal{X}}$ is a boundary point of $\mathcal{P}$} if $\mathbf{x}\not\in{\mathcal{P}}^{\circ}$ and $\mathbf{x}\not\in{\mathcal{P}}^{\perp}$. The boundary of $\mathcal{P}$}, written $\partial{\mathcal{P}}$, is the set of all boundary points of $\mathcal{P}$}.
A subset $\mathcal{P}$ in $\cal X$} is regular open if it coincides with the interior of its closure. For $\mathcal{X}=\mathbb{R}^{2}$, a subset $\mathcal{S}\subset\mathbb{R}^{2}$ is semianalytic if there exist a finite number of analytic functions $g_{i}:\mathbb{R}^{2}\rightarrow\mathbb{R}$ such that $\mathcal{S}$ is in the universe of a finite Boolean algebra formed from the sets $\mathcal{X}_{i}=\left\{\mathbf{x}\in\mathbb{R}^{2}:g_{i}(\mathbf{x})\geq 0\right\}$. In particular, a semianalytic set is semialgebraic if all the $g_{i}$’s are polynomials. These concepts lead to
Definition 2.1 (Yin Space [40]).
A Yin set $\mathcal{Y}\subseteq\mathbb{R}^{2}$ is a regular open semianalytic set with bounded boundary. The Yin space $\mathbb{Y}$ is the class of all Yin sets.
In Definition 2.1, the regularity captures the salient feature that continua are free of low-dimensional elements such as isolated points and crevices, the openness leads to a unique boundary representation of any Yin set, and the semianalyticity ensures that a finite number of entities suffice for the boundary representation.
Each Yin set $\mathcal{Y}\neq\emptyset,\mathbb{R}^{2}$ can be uniquely expressed [40, Corollary 3.13] as
(3) $\mathcal{Y}=\cup^{\perp\perp}_{j}{\cal Y}_{j}=\cup^{\perp\perp}_{j}\cap_{i}\operatorname{int}(\gamma_{j,i}),$
where ${\cal Y}_{j}$ is the $j$th connected component of $\mathcal{Y}$, the binary operation $\cup^{\perp\perp}$ the regularized union defined as $\mathcal{S}\cup^{\perp\perp}\mathcal{T}:=\left(\mathcal{S}\cup\mathcal{T}\right)^{\perp\perp}$, $\{\gamma_{j,i}\}$ the set of pairwise almost disjoint Jordan curves satisfying $\partial{\cal Y}_{j}=\cup_{i}\gamma_{j,i}$, and $\operatorname{int}(\gamma_{j,i})$ the complement of $\gamma_{j,i}$ that always lies at the left of an observer who traverses $\gamma_{j,i}$ according to its orientation.
Theorem 2.2 (Zhang and Li [40]).
$\left(\mathbb{Y},\cup^{\perp\perp},\cap,^{\perp},\emptyset,\mathbb{R}^{2}\right)$ is a Boolean algebra.
Corollary 2.3.
Denote by $\mathbb{Y}_{\mathrm{c}}$ the subspace of $\mathbb{Y}$ where the boundary of each Yin set is constituted by a finite number of cubic splines. Then $\mathbb{Y}_{\mathrm{c}}$ is a sub-algebra of $\mathbb{Y}$.
The above Boolean algebra is implemented in [40]. In this work, the problem domain $\Omega$ in (1) is assumed to be in $\mathbb{Y}$ and approximated by a Yin set in $\mathbb{Y}_{\mathrm{c}}$.
3 Partitioning a Yin set into cut cells
The arbitrary complexity of $\Omega$ is handled by a divide-and-conquer approach: in Section 3.1 we cut $\Omega$ by a Cartesian grid to generate a set $\mathsf{C}_{\Omega}$ of cut cells. In Section 3.2, we merge adjacent cut cells so that, for a user-specified value $\epsilon\in(0,1)$, the volume fraction of each cut cell is no less than $\epsilon$, thus preventing small volumes of the cut cells to ensure good conditioning of spatial discretizations in Section 4.
3.1 Cut cells
We embed the domain $\Omega$ inside an open rectangle $\Omega_{R}\supset\Omega$ and divide $\Omega_{R}$ by a Cartesian grid into square control volumes or cells,
(4) $\mathbf{C}_{\mathbf{i}}\coloneqq\left(\mathbf{i}h,(\mathbf{i}+\mathbbm{1})h\right),$
where $h$ is the uniform grid size, $\mathbf{i}\in\mathbb{Z}^{2}$ a multi-index, and $\mathbbm{1}\in\mathbb{Z}^{2}$ the multi-index with all components equal to one. Note that the assumption of $h$ being uniform is for ease of exposition only, as our algorithm also applies to non-uniform grids.
We initialize the $\mathbf{i}$th cut cell as $\mathcal{C}_{\mathbf{i}}=\mathbf{C}_{\mathbf{i}}\cap\Omega$ and call $\mathcal{C}_{\mathbf{i}}$ an empty cell if $\mathcal{C}_{\mathbf{i}}=\emptyset$, a regular cell if $\mathcal{C}_{\mathbf{i}}=\mathbf{C}_{\mathbf{i}}$, or an irregular cell otherwise. Along the $d$th dimension, the higher face and the lower face of the $\mathbf{i}$th cell $\mathbf{C}_{\mathbf{i}}$ are respectively given by
(5) $\mathrm{F}_{\mathbf{i}+\frac{1}{2}\bm{\mathrm{e}}^{d}}:=\left((\mathbf{i}+\bm{\mathrm{e}}^{d})h,(\mathbf{i}+\mathbbm{1})h\right),\quad\mathrm{F}_{\mathbf{i}-\frac{1}{2}\bm{\mathrm{e}}^{d}}:=\left(\mathbf{i}h,(\mathbf{i}+\mathbbm{1}-\bm{\mathrm{e}}^{d})h\right),$
where $\bm{\mathrm{e}}^{d}\in\mathbb{Z}^{2}$ is the multi-index whose components are all zero except the $d$th component being one. The higher/lower cut faces and the cut boundary of the $\mathbf{i}$th cell $\mathbf{C}_{\mathbf{i}}$ are respectively given by
(6) $\mathcal{F}_{\mathbf{i}\pm\frac{1}{2}\bm{\mathrm{e}}^{d}}:=\mathrm{F}_{\mathbf{i}\pm\frac{1}{2}\bm{\mathrm{e}}^{d}}\cap\Omega\quad\text{ and }\quad\mathcal{B}_{\mathbf{i}}:=\mathbf{C}_{\mathbf{i}}\cap\partial\Omega.$
For a domain $\Omega$ and its embedding rectangle $\Omega_{R}$, the set of cut cells is defined as
(7) $\mathsf{C}_{\Omega}:=\left\{\mathcal{C}_{\mathbf{i}}:\mathcal{C}_{\mathbf{i}}\neq\emptyset;\ \cup_{\mathbf{i}}^{\bot\bot}\mathcal{C}_{\mathbf{i}}=\Omega\right\}.$
Thanks to the rectangular structure of the Cartesian grid, the set of neighbors of a cut cell $\mathcal{C}_{\mathbf{j}}$ is easily obtained as $\mathsf{N}_{\mathbf{j}}:=\left\{\mathcal{C}_{\mathbf{i}}:\mathcal{C}_{\mathbf{i}}\in\mathsf{C}_{\Omega},\ \|\mathbf{i}-\mathbf{j}\|_{1}=1\right\}$. The connected components of $\mathcal{C}_{\mathbf{j}}$ are denoted by $\mathcal{C}_{\mathbf{j}}^{1}$, $\mathcal{C}_{\mathbf{j}}^{2}$, $\ldots$ so that $\mathcal{C}_{\mathbf{j}}=\cup^{\perp\perp}_{k}\mathcal{C}_{\mathbf{j}}^{k}$. The set of neighboring components of a cut-cell component $\mathcal{C}_{\mathbf{j}}^{k}$ is defined as
(8) $\mathsf{N}_{\mathbf{j}}^{k}:=\left\{\mathcal{C}_{\mathbf{i}}^{\ell}:\mathcal{C}_{\mathbf{i}}\in\mathsf{C}_{\Omega},\ \|\mathbf{i}-\mathbf{j}\|_{1}=1,\ \overline{\mathcal{C}_{\mathbf{j}}^{k}}\cap\overline{\mathcal{C}_{\mathbf{i}}^{\ell}}\neq\emptyset\right\}.$
Note that a neighboring cut cell $\mathcal{C}_{\mathbf{i}}$ might have multiple components in $\mathsf{N}_{\mathbf{j}}^{k}$.
3.2 Resolving the small-volume problem by cell merging
In practice, the volume $\lVert\mathcal{C}_{\mathbf{i}}\rVert$ of a cut cell $\mathcal{C}_{\mathbf{i}}$ may be very small, leading to ill-conditioning of the discretized operator. This problem has been addressed by a number of approaches such as cell merging [16], redistribution [1], and special discretization schemes [10].
Our solution of the small-volume problem is a novel cell-merging algorithm whose output is $\mathsf{C}_{\epsilon}^{h}(\Omega)$, a regularized set of cut cells of $\Omega$ where each cut cell has only one connected component and the volume fraction of this component is no less than a user-specified lower bound $\epsilon$:
(9) $\mathsf{C}_{\epsilon}^{h}(\Omega):=\left\{\mathcal{C}_{\mathbf{i}}:\mathcal{C}_{\mathbf{i}}=\mathcal{C}_{\mathbf{i}}^{1}\neq\emptyset;\ \cup_{\mathbf{i}}^{\bot\bot}\mathcal{C}_{\mathbf{i}}=\Omega;\ \|{\cal C}_{\mathbf{i}}\|\geq\epsilon h^{2}\right\},$
where $\mathcal{C}_{\mathbf{i}}=\mathcal{C}_{\mathbf{i}}^{1}$ means that each $\mathcal{C}_{\mathbf{i}}$ consists of only one connected component. As a design choice to ensure that the unknown function $u$ in (1) has only one cell average per cut cell, the condition $\mathcal{C}_{\mathbf{i}}=\mathcal{C}_{\mathbf{i}}^{1}$ retains the simplicity of rectangular grids and facilitates the design of the multigrid solver in Section 5.
Algorithm 1 CutAndMergeCells
Input: $\Omega\in\mathbb{Y}_{c}$: the problem domain;
$\parbox[t]{35.56593pt}{}\Omega_{R}$: the rectangle that contains $\Omega$;
$\parbox[t]{35.56593pt}{}h$: the size of the Cartesian grid that discretizes $\Omega_{R}$;
$\parbox[t]{35.56593pt}{}\epsilon\in(0,1)$: a user-specified threshold of small volume fractions.
Precondition: $h$ is sufficiently small to resolve the topology and geometry of $\Omega$.
Output: $\mathsf{C}_{\epsilon}^{h}$: the set of cut cells of $\Omega$ in (9).
[…]
2:for each cut cell $\mathcal{C}_{\mathbf{i}}=\left(\cup^{\bot\bot}\mathcal{C}_{\mathbf{i}}^{k}\right)_{k=1}^{n_{\mathbf{i}}}$ with $n_{\mathbf{i}}\geq 2$ connected components do
3:  $\mathcal{C}_{\mathbf{i}}\leftarrow\mathcal{C}_{\mathbf{i}}^{m}$ where $m=\operatorname*{argmax}_{k=1}^{n_{\mathbf{i}}}\{\|\mathcal{C}_{\mathbf{i}}^{k}\|\}$
4:  for each $k=1,\ldots,m-1,m+1,\ldots,n_{\mathbf{i}}$ do
5:   $\mathcal{C}_{\mathbf{j}}^{i}\leftarrow\mathcal{C}_{\mathbf{j}}^{i}\cup^{\bot\bot}\mathcal{C}_{\mathbf{i}}^{k}$ where $\mathcal{C}_{\mathbf{j}}^{i}=\arg\min_{{\cal C}_{\mathbf{j}^{\prime}}^{\ell}\in\mathsf{N}_{\mathbf{i}}^{k}}\left|\|\mathcal{C}_{\mathbf{j}^{\prime}}^{\ell}\|+\|\mathcal{C}_{\mathbf{i}}^{k}\|-h^{2}\right|$
6:  end for
7:end for
8:for each cut cell $\mathcal{C}_{\mathbf{i}}$ satisfying $\|\mathcal{C}_{\mathbf{i}}\|<\epsilon h^{2}$ do
9:  $\mathcal{C}_{\mathbf{j}}\leftarrow\mathcal{C}_{\mathbf{j}}\cup^{\bot\bot}\mathcal{C}_{\mathbf{i}}$ where $\mathcal{C}_{\mathbf{j}}=\arg\min_{{\cal C}_{\mathbf{j}^{\prime}}\in\mathsf{N}_{\mathbf{i}}^{1}}\left|\|\mathcal{C}_{\mathbf{j}^{\prime}}\|+\|\mathcal{C}_{\mathbf{i}}\|-h^{2}\right|$
10:  $\mathsf{C}_{\epsilon}^{h}\leftarrow\mathsf{C}_{\epsilon}^{h}\setminus\mathcal{C}_{\mathbf{i}}$
11:end for
12:return $\mathsf{C}_{\epsilon}^{h}$
[…]
[width=0.4]tikz/Cutcell
Figure 1: An illustration of Algorithm 1 in generating $\mathsf{C}_{\epsilon}^{h}(\Omega)$ in (9) with $\epsilon=0.2$ by cutting and merging cells for the domain $\Omega$. The cut cells ${\cal C}_{\mathbf{l}}$, ${\cal C}_{\mathbf{p}}$, and ${\cal C}_{\mathbf{q}}$ are regular, empty, and irregular, respectively. The symbol $\leftrightarrow$” indicates cell merging. Originally, ${\cal C}_{\mathbf{k}}\in\mathsf{C}_{\Omega}$ is an irregular cell with two small connected components, which are merged to $\mathcal{C}_{\mathbf{j}}$ at line 5 and $\mathcal{C}_{\mathbf{i}}$ at line 9, respectively. Then ${\cal C}_{\mathbf{k}}$ is removed from $\mathsf{C}_{\epsilon}^{h}(\Omega)$ and its type changed from “irregular” to “empty.” The type of ${\cal C}_{\mathbf{i}}$ is changed from “regular” to “irregular” whereas the type of ${\cal C}_{\mathbf{j}}$ remains “irregular.”
In Algorithm 1, we first initialize $\mathsf{C}_{\epsilon}^{h}$ with $\mathsf{C}_{\Omega}$ in (7) in line 1. For each cut cell ${\cal C}_{\mathbf{i}}\in\mathsf{C}_{\epsilon}^{h}$ with multiple components, we identify ${\cal C}_{\mathbf{i}}^{m}$ as a component of ${\cal C}_{\mathbf{i}}$ with the maximum volume, set ${\cal C}_{\mathbf{i}}={\cal C}_{\mathbf{i}}^{m}$, and merge any other component to a neighboring component ${\cal C}_{\mathbf{j}}^{i}$ so that the volume fraction of the merged component is closest to 1; see line 5 and (8). When the grid size $h$ is too large, the geometry of the boundary may not be well resolved so that at the exit of the double loop in lines 2–7 there may still exist multiple cut cells associated with a single control volume. However, so long as $h$ is sufficiently small, the loop in lines 2–7 would leave each cut cell with only one component. Finally in lines 8–11, each small cut cell ${\cal C}_{\mathbf{i}}$ is merged to a neighboring cell and removed from the set of cut cells.
An example of Algorithm 1 is shown in Figure 1. After line 1, the cut cell $\mathcal{C}_{\mathbf{k}}\in\mathsf{C}_{\Omega}$ has two connected components: $\mathcal{C}_{\mathbf{k}}^{1}$ on the left has a larger volume than $\mathcal{C}_{\mathbf{k}}^{2}$ on the right. During the double loop in lines 2–7, $\mathcal{C}_{\mathbf{k}}$ is set to $\mathcal{C}_{\mathbf{k}}^{1}$ and $\mathcal{C}_{\mathbf{k}}^{2}$ is merged with $\mathcal{C}_{\mathbf{j}}^{1}$. During the loop in lines 8–11, all cut cells with small volume fractions are merged with a neighboring cut cell; in particular, $\mathcal{C}_{\mathbf{k}}$ is merged with $\mathcal{C}_{\mathbf{i}}$.
4 Discretizing equation (1) into a linear system of cell averages
The cell average of a scalar function $\phi:\overline{\Omega}\rightarrow\mathbb{R}$ over a cut cell $\mathcal{C}_{\mathbf{i}}\in\mathsf{C}_{\epsilon}^{h}(\Omega)$ is defined as
(10) $\left\langle\phi\right\rangle_{\mathbf{i}}:=\frac{1}{\lVert\mathcal{C}_{\mathbf{i}}\rVert}\int_{\mathcal{C}_{\mathbf{i}}}\phi(\mathbf{x})\ \mathrm{d}\mathbf{x},$
where $\lVert\mathcal{C}_{\mathbf{i}}\rVert$ is the volume of $\mathcal{C}_{\mathbf{i}}$; similarly, the face average over the cut face $\mathcal{F}_{\mathbf{i}+\frac{1}{2}\bm{\mathrm{e}}^{d}}$ and the boundary average over the cut boundary $\mathcal{B}_{\mathbf{i}}$ are respectively
[…]
where $\|\cdot\|$ denotes the length of a cut face or cut boundary in (6).
[…]
4.1 Discretizing (1) on SFV cells
A face $\mathrm{F}_{\mathbf{i}+\frac{1}{2}\bm{\mathrm{e}}^{d}}$ or $\mathrm{F}_{\mathbf{i}-\frac{1}{2}\bm{\mathrm{e}}^{d}}$ in (5) is said to be extendable if it is entirely contained in $\partial\Omega$. Write
(12) ${\cal S}_{\mathbf{i}}^{d,+}:=\{{\cal C}_{\mathbf{i}-m\bm{\mathrm{e}}^{d}}:m=0,1,2,3\},\quad{\cal S}_{\mathbf{i}}^{d,-}:=\{{\cal C}_{\mathbf{i}+m\bm{\mathrm{e}}^{d}}:m=0,1,2,3\}.$
A cut cell $\mathcal{C}_{\mathbf{i}}\in\mathsf{C}_{\epsilon}^{h}(\Omega)$ is extendable in the high direction along the $d$th dimension if the face $\mathrm{F}_{\mathbf{i}+\frac{1}{2}\bm{\mathrm{e}}^{d}}$ is extendable and all cut cells in ${\cal S}_{\mathbf{i}}^{d,+}$ are regular; similarly, $\mathcal{C}_{\mathbf{i}}$ is extendable in the low direction along the $d$th dimension if the face $\mathrm{F}_{\mathbf{i}-\frac{1}{2}\bm{\mathrm{e}}^{d}}$ is extendable and all cut cells in ${\cal S}_{\mathbf{i}}^{d,-}$ are regular.
Figure 2: An example of ghost filling near the regular boundary. $\mathrm{F}_{\mathbf{i}+\frac{1}{2}\bm{\mathrm{e}}^{d}}$ is an extendable face and ${\cal C}_{\mathbf{i}}$ is an extendable cell in the high direction along the first dimension.
For an extendable cell, we append two (regular) ghost cells to each extendable face in the corresponding direction and smoothly extend cell averages of $\phi$ to the ghost cells. For the example in Figure 2, we use
(13) $\begin{array}[]{l}\left\langle\phi\right\rangle_{\mathbf{i}+\mathbf{e}^{d}}=\frac{1}{12}\left(3\left\langle\phi\right\rangle_{\mathbf{i}-3\mathbf{e}^{d}}-17\left\langle\phi\right\rangle_{\mathbf{i}-2\mathbf{e}^{d}}+43\left\langle\phi\right\rangle_{\mathbf{i}-\mathbf{e}^{d}}-77\left\langle\phi\right\rangle_{\mathbf{i}}+60\left\langle\phi\right\rangle_{\mathbf{i}+\frac{1}{2}\mathbf{e}^{d}}\right)+O(h^{5});\\ \left\langle\phi\right\rangle_{\mathbf{i}+2\mathbf{e}^{d}}=\frac{1}{12}\left(27\left\langle\phi\right\rangle_{\mathbf{i}-3\mathbf{e}^{d}}-145\left\langle\phi\right\rangle_{\mathbf{i}-2\mathbf{e}^{d}}+335\left\langle\phi\right\rangle_{\mathbf{i}-\mathbf{e}^{d}}-505\left\langle\phi\right\rangle_{\mathbf{i}}+75\left\langle\phi\right\rangle_{\mathbf{i}+\frac{1}{2}\mathbf{e}^{d}}\right)+O(h^{5})\end{array}$
to fill ghost cells while enforcing the Dirichlet boundary condition $\left\langle\phi\right\rangle_{\mathbf{i}+\frac{1}{2}\mathbf{e}^{d}}$. As for the Neumann boundary condition $\left\langle\frac{\partial\phi}{\partial\mathbf{n}}\right\rangle_{\mathbf{i}+\frac{1}{2}\mathbf{e}^{d}}$, the ghost-filling formulas are
(14) $\begin{array}[]{l}\left\langle\phi\right\rangle_{\mathbf{i}+\mathbf{e}^{d}}=\frac{1}{10}\left(\left\langle\phi\right\rangle_{\mathbf{i}-3\mathbf{e}^{d}}-5\left\langle\phi\right\rangle_{\mathbf{i}-2\mathbf{e}^{d}}+9\left\langle\phi\right\rangle_{\mathbf{i}-\mathbf{e}^{d}}+5\left\langle\phi\right\rangle_{\mathbf{i}}+12h\left\langle\frac{\partial\phi}{\partial\mathbf{n}}\right\rangle_{\mathbf{i}+\frac{1}{2}\mathbf{e}^{d}}\right)+O(h^{5}),\\ \left\langle\phi\right\rangle_{\mathbf{i}+2\mathbf{e}^{d}}=\frac{1}{2}\left(3\left\langle\phi\right\rangle_{\mathbf{i}-3\mathbf{e}^{d}}-15\left\langle\phi\right\rangle_{\mathbf{i}-2\mathbf{e}^{d}}+29\left\langle\phi\right\rangle_{\mathbf{i}-\mathbf{e}^{d}}-15\left\langle\phi\right\rangle_{\mathbf{i}}+12h\left\langle\frac{\partial\phi}{\partial\mathbf{n}}\right\rangle_{\mathbf{i}+\frac{1}{2}\mathbf{e}^{d}}\right)+O(h^{5}).\end{array}$
For periodic boundary conditions, the values of ghost cells are copied directly from those of the corresponding regular cells inside the domain $\Omega$.
Definition 4.1.
Recall ${\cal L}=a\frac{\partial^{2}}{\partial x^{2}}+b\frac{\partial^{2}}{\partial x\partial y}+c\frac{\partial^{2}}{\partial y^{2}}$ from (1) and write
(15) $\mathcal{S}_{\mathbf{i}}:=\begin{cases}\{{\cal C}_{\mathbf{j}}:\mathbf{j}=\mathbf{i}+m\bm{\mathrm{e}}^{d};d=1,2;m=0,\pm 1,\pm 2\}&\text{if }b=0;\\ \{{\cal C}_{\mathbf{j}}:\mathbf{j}=\mathbf{i}+m_{1}\bm{\mathrm{e}}^{1}+m_{2}\bm{\mathrm{e}}^{2};m_{1},m_{2}=0,\pm 1,\pm 2\}&\text{if }b\neq 0.\end{cases}$
A cut cell $\mathcal{C}_{\mathbf{i}}\in\mathsf{C}_{\epsilon}^{h}(\Omega)$ is called an SFV cell if each cut cell in $\mathcal{S}_{\mathbf{i}}$ is either a regular cell or a ghost cell; otherwise it is called a PLG cell.
The case $b=0$ in (15) follows directly from (12) and the following symmetric finite-volume discretization of the first and second derivatives:
(16a) $\left\langle\frac{\partial\phi}{\partial x_{d}}\right\rangle_{\mathbf{i}}=\frac{1}{12h}\Big(\left\langle\phi\right\rangle_{\mathbf{i}-2\bm{\mathrm{e}}^{d}}-8\left\langle\phi\right\rangle_{\mathbf{i}-\bm{\mathrm{e}}^{d}}+8\left\langle\phi\right\rangle_{\mathbf{i}+\bm{\mathrm{e}}^{d}}-\left\langle\phi\right\rangle_{\mathbf{i}+2\bm{\mathrm{e}}^{d}}\Big)+\operatorname{O}(h^{4}),$
(16b) $\left\langle\frac{\partial^{2}\phi}{\partial x_{d}^{2}}\right\rangle_{\mathbf{i}}=\frac{1}{12h^{2}}\Big(-\left\langle\phi\right\rangle_{\mathbf{i}-2\bm{\mathrm{e}}^{d}}+16\left\langle\phi\right\rangle_{\mathbf{i}-\bm{\mathrm{e}}^{d}}-30\left\langle\phi\right\rangle_{\mathbf{i}}+16\left\langle\phi\right\rangle_{\mathbf{i}+\bm{\mathrm{e}}^{d}}-\left\langle\phi\right\rangle_{\mathbf{i}+2\bm{\mathrm{e}}^{d}}\Big)+\operatorname{O}(h^{4});$
see [39] for a proof of the fourth-order accuracy. The case $b\neq 0$ in (15) follows from the discretization of the cross derivative $\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}(i\neq j)$ by applying (16a) first in the $i$th direction and then in the $j$th direction.
4.2 Discretizing (1) on PLG cells
In Section 4.2.1, we briefly review the PLG algorithm [41] that generates a suitable stencil for each PLG cell. In Section 4.2.2, we fit a multivariate polynomial locally to discretize $\left\langle{\cal L}u\right\rangle_{\mathbf{i}}$ as a linear combination of cell averages and boundary averages.
4.2.1 Poised lattice generation (PLG)
We start with
Definition 4.2 (Lagrange interpolation problem (LIP)).
Denote by $\Pi_{n}^{{\scriptsize\textsf{D}}}$ the vector space of all D-variate polynomials of degree no more than $n$ with real coefficients. Given points $\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{N}\in\mathbb{R}^{\scriptsize\textsf{D}}$ and the same number of data $f_{1},f_{2},\cdots,f_{N}\in\mathbb{R}$, the Lagrange interpolation problem seeks a polynomial $f\in\Pi_{n}^{\scriptsize\textsf{D}}$ such that
(17) $\forall j=1,2,\cdots,N,\quad f(\mathbf{x}_{j})=f_{j},$
where $\Pi_{n}^{\scriptsize\textsf{D}}$ is the interpolation space and $\{\mathbf{x}_{j}\}_{j=1}^{N}$ are the interpolation sites.
The sites $\left\{\mathbf{x}_{j}\right\}_{j=1}^{N}$ are poised in $\Pi_{n}^{\scriptsize\textsf{D}}$ if, for any given data $\left\{f_{j}\right\}_{j=1}^{N}$, there exists a unique $f\in\Pi_{n}^{\scriptsize\textsf{D}}$ satisfying (17). For a basis $\left\{\phi_{j}\right\}$ of $\Pi_{n}^{\scriptsize\textsf{D}}$, $\left\{\mathbf{x}_{j}\right\}_{j=1}^{N}$ are poised if and only if $N=\dim\Pi_{n}^{\scriptsize\textsf{D}}=\binom{n+{\scriptsize\textsf{D}}}{n}$ and the following sample matrix $M\in\mathbb{R}^{N\times N}$ is non-singular,
(18) $\forall j,k=1,2,\ldots N,\quad M(\left\{\phi_{j}\right\};\left\{\mathbf{x}_{k}\right\})=\Big[M_{jk}\Big]:=\Big[\phi_{j}(\mathbf{x}_{k})\Big].$
For ${\scriptsize\textsf{D}}=1$, the LIP is unisolvent if and only if its sites are pairwise distinct. For ${\scriptsize\textsf{D}}>1$, however, it is difficult to determine whether a set of sites is poised in $\Pi^{{\scriptsize\textsf{D}}}_{n}$. For example, the lattice $\{(5,0),(-5,0),(0,5),(0,-5),(4,3),(-3,4)\}$ is not poised in $\Pi^{2}_{2}=\mathrm{span}(1,x,y,x^{2},y^{2},xy)$ because the corresponding sample matrix in (18) is singular. As the core difficulty of multivariate interpolation, the poisedness of interpolation sites in multiple dimensions depends on their geometric configuration.
Definition 4.3 (PLG in $\mathbb{Z}^{\scriptsize\textsf{D}}$ [41]).
Given a finite set $K\subset\mathbb{Z}^{\text{D}}$ of feasible nodes, a starting point $\mathbf{q}\in K$, and a degree $n\in\mathbb{Z}^{+}$, the problem of poised lattice generation is to choose a lattice $\mathcal{T}\subset K$ such that $\mathcal{T}$ is poised in $\Pi_{n}^{\scriptsize\textsf{D}}$ and $\mathcal{T}=\dim\Pi_{n}^{\scriptsize\textsf{D}}$.
[…]
Via a fusion of approximation theory, group theory, and search algorithms in artificial intelligence, we solved the PLG problem in Definition 4.3 by a novel and efficient PLG algorithm [41], which is applied in this work to discretize $\left\langle{\cal L}u\right\rangle_{\mathbf{i}}$ with $K$.
4.2.2 Approximating $\left\langle{\cal L}u\right\rangle_{\mathbf{i}}$ with a linear combination of cell averages (and a boundary average)
Let $\mathcal{S}_{\mathrm{PLG}}(\mathbf{i})=\left\{\mathcal{C}_{\mathbf{j}_{1}},\mathcal{C}_{\mathbf{j}_{2}},\cdots,\mathcal{C}_{\mathbf{j}_{N}}\right\}$ denote the poised lattice generated by the PLG algorithm where $N=\dim{\Pi_{n}^{\scriptsize\textsf{D}}}$. As shown in Figure 3, the stencil for discretizing $\mathcal{L}$ over a PLG cell $\mathcal{C}_{\mathbf{i}}$ is
[…]
[width=0.36]tikz/Stencil
Figure 3: An example of the stencil for multivariate polynomial fitting in the FV formulation for ${\scriptsize\textsf{D}}=2$ and $n=4$. $\bullet$” marks the starting point $\mathbf{q}=\mathbf{i}$, the cells with dark shades constitute $\mathcal{S}_{\mathrm{PLG}}$ in (19), and the thick curve segment represents the cut boundary $\mathcal{B}_{\mathbf{i}}$.
Given the cell averages and the boundary average
[…]
the goal is to determine a vector of coefficients $\bm{\beta}=\left[\beta_{1},\cdots,\beta_{N},\beta_{\mathrm{b}}\right]^{\top}$ such that the linear combination $\bm{\beta}^{\top}\overline{\mathbf{u}}$ is an (n1)(n-1)th-order approximation of $\left\langle{\cal L}u\right\rangle_{\mathbf{i}}$,
(21) $\forall u\in{\cal C}^{n+1}(\mathbb{R}^{\scriptsize\textsf{D}},\mathbb{R}),\quad\bm{\beta}^{\top}\overline{\mathbf{u}}=\left\langle\mathcal{L}u\right\rangle_{\mathbf{i}}+\operatorname{O}(h^{n-1}),$
[…]
The equations on $\bm{\beta}$ are obtained via a restricted version of (21),
[…]
(23) $M\bm{\beta}=\overline{\bm{\phi}},$
where $\overline{\bm{\phi}}=\left(\left\langle\mathcal{L}\phi_{1}\right\rangle_{\mathbf{i}},\left\langle\mathcal{L}\phi_{2}\right\rangle_{\mathbf{i}},\ldots,\left\langle\mathcal{L}\phi_{N}\right\rangle_{\mathbf{i}}\right)^{\top}\in\mathbb{R}^{N}$; for an irregular PLG cell ${\cal C}_{\mathbf{i}}$, we have
[…]
For a regular PLG cell ${\cal C}_{\mathbf{i}}$, the last column of $M$ in (24) is dropped, so are the last elements of $\bm{\beta}$ and $\overline{\bm{\phi}}$.
We calculate the integrals on regular cells by six-order recursive Gauss formulas. In contrast, the integral of a scalar function over an irregular cut cell is first converted by Green’s theorem to another integral on the boundary Jordan curve and then approximated by sixth-order Gauss formulas; see [36] for more details. Together with the explicit approximation of the boundary Jordan curve with cubic splines, this integral formulation makes our method robust in that its fourth accuracy holds even at the presence of kinks on the domain boundary.
If (23) is under-determined, we solve a constrained optimization problem
(25) $\min_{\bm{\beta}\in\mathbb{R}^{N+1}}\left\|\bm{\beta}\right\|_{W^{-1}}\quad\mathrm{s.t.}\quad M\bm{\beta}=\overline{\bm{\phi}},$
where the square matrix $W$ is symmetric positive definite, the $W$-inner product of two vectors is $\langle\mathbf{w},\mathbf{v}\rangle_{W}:=\mathbf{w}^{\top}W\mathbf{v}$, and the $W$-norm of a vector is $\|\mathbf{v}\|_{W}:=\sqrt{\langle\mathbf{v},\mathbf{v}\rangle_{W}}$.
Since $M$ has full row rank, it follows from Lemma 4.4 that (25) is solved by
(26) $\bm{\beta}_{\min}=WM^{\top}\left(MWM^{\top}\right)^{-1}\overline{\bm{\phi}}.$
Lemma 4.4.
Let $A\in\mathbb{R}^{m\times N}$ be a matrix with full column rank and $W\in\mathbb{R}^{m\times m}$ be a symmetric positive definite matrix. For any $\mathbf{v}\in\mathbb{R}^{N}$, the optimization problem of $\min_{\mathbf{x}\in\mathbb{R}^{m}}\left\|\mathbf{x}\right\|_{W^{-1}}$ under the constraint $A^{\top}\mathbf{x}=\mathbf{v}$ admits a unique solution $\mathbf{x}_{\min}=WA\left(A^{\top}WA\right)^{-1}\mathbf{v}$.
Proof 4.5.
See [12, §5.3, §5.6, §6.1].
It is reasonable to demand that a cut cell closer to ${\cal C}_{\mathbf{i}}$ has a greater influence upon the linear system than a cut cell away from ${\cal C}_{\mathbf{i}}$. To this end, we set
(27a) $W^{-1}$ $=\operatorname{diag}\left(w_{1},\cdots,w_{N},w_{\mathrm{b}}\right);$
(27b) $w_{k}$ $=\max\left\{\|\mathbf{j}_{k}-\mathbf{i}\|_{2},w_{\min}\right\};\quad w_{\mathrm{b}}=w_{\min}=0.5,$
where the nonzero value of $w_{\min}$ prevents $\frac{1}{w_{k}}$ from being too large.
To maintain good conditioning, the basis $\left\{\phi_{j}\right\}_{j=1}^{N}$ of $\Pi_{n}^{\scriptsize\textsf{D}}$ is set to
(28) $\begin{array}[]{l}\Phi_{n}^{\scriptsize\textsf{D}}(h;\mathbf{p})=\left\{\left(\frac{\mathbf{x}-\mathbf{p}}{h}\right)^{\bm{\alpha}}:\bm{\alpha}\in\{0,1,\ldots,n\}^{{\scriptsize\textsf{D}}}\text{ and }\par\|\bm{\alpha}\|_{1}\in[0,n]\right\},\end{array}$
where $\mathbf{p}\in\mathbb{R}^{\scriptsize\textsf{D}}$ is the center of ${\cal C}_{\mathbf{i}}$.
4.3 The linear system as the discrete elliptic problems
Given $\Omega$, $\Omega_{R}$, $h$, and $\epsilon$, Algorithm 1 uniquely determines the set $\mathsf{C}^{h}_{\epsilon}$ of cut cells in (9). For SFV cells, the symmetric FV formulas in Section 4.1 are employed to discretize the integral of (1). For each PLG cell ${\cal C}_{\mathbf{i}}$, the vector $\overline{\bm{\phi}}$ and the matrices $M$ and $W$ in (23) and (27) yield $\bm{\beta}_{\min}$ in (26), which, together with (21), implies that $\bm{\beta}_{\min}^{\top}\overline{\mathbf{u}}$ is an (n1)(n-1)th-order approximation of the integral $\left\langle\mathcal{L}u\right\rangle_{\mathbf{i}}$ of (1) over ${\cal C}_{\mathbf{i}}$. Combine the two cases and we have a linear system of the form
(29) $A\mathbf{u}=\mathbf{b}:=\mathbf{f}-N\mathbf{g},$
[…]
(30) $\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{1}\\ \mathbf{u}_{2}\end{bmatrix}=\begin{bmatrix}\mathbf{b}_{1}\\ \mathbf{b}_{2}\end{bmatrix}$,
where $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$ contain cell averages of $u$ over SFV cells and PLG cells, respectively. The eigenvalues of $A_{11}$ have nonnegative real parts. In contrast, each of $A_{12}$, $A_{21}$, and $A_{22}$ is asymmetric and indefinite; all we know about them is their sparsity.
The error and the residual of an approximate solution $\tilde{\mathbf{u}}\approx\mathbf{u}$ of (29) are respectively defined as
(31) $\mathbf{e}(\tilde{\mathbf{u}}):=\mathbf{u}-\tilde{\mathbf{u}},\quad\mathbf{r}(\tilde{\mathbf{u}}):=\mathbf{b}-A\tilde{\mathbf{u}}$.
Then (29) can be rewritten as the equivalent residual equation $A\mathbf{e}=\mathbf{r}=\mathbf{b}-A\tilde{\mathbf{u}}$, which is conducive to the design and exposition of multigrid methods.
5 The cut-cell geometric multigrid method
The PLG discretization results in the indefiniteness of the block matrix $A_{22}$ in (30), and thus prohibits a direct application of traditional geometric multigrid methods. To cope with this difficulty, we give a total ordering to the set of PLG cells and prove in Section 5.1 that the LU factorization of the corresponding subblock $A_{22}$ has the optimal complexity of $O(h^{-1})$. Then we design a fixed-point iteration in Section 5.2 as a block smoother of (30) and assemble these components into a cut-cell V-cycle in Section 5.3. In Section 5.4, we assemble these components to propose a cut-cell full multigrid (FMG) cycle as a new cut-cell geometric multigrid method that solves the block linear system (30) with the optimal complexity of $O(h^{-2})$.
5.1 An optimal LU factorization of $A_{22}$ in (30)
The bandwidth of a square matrix $A$ is the minimum nonnegative integer $k$ such that |ij|>k|i-j|>k implies $a_{i,j}=0$. The bandwidth of $A_{22}$ in (30) is greatly affected by the ordering of the unknowns in $\mathbf{u}_{2}$, i.e., the ordering of the PLG cells. By the unique representation of Yin sets in (3), it suffices to define the ordering for PLG cells close to a single Jordan curve $\gamma:[0,1]\to\mathbb{R}^{2}$ where $\gamma(0)=\gamma(1)$ and $\gamma|_{[0,1)}$ is a continuous injection.
Definition 5.1.
For a Jordan curve $\gamma\subset\partial\Omega$, the total ordering on the multi-index set $\mathcal{I}_{\mathrm{PLG},\gamma}:=\{\mathbf{i}:{\cal C}_{\mathbf{i}}\text{ is a PLG cell near }\gamma\}$ is given by $\mathbf{i}\leq\mathbf{j}$ if and only if $s(\mathbf{i})\leq s(\mathbf{j})$, where $s(\mathbf{i})$ is the parameter of the point $\gamma(s(\mathbf{i}))$ on $\gamma$ that is closest to the center of the cell $\mathbf{C}_{\mathbf{i}}$ in (4).
[…]
[width=0.5]tikz/OrderOnBdry
Figure 4: Illustrating the total ordering of PLG cells in Definition 5.1. The SFV and PLG cells are shaded in gray and in yellow, respectively. The dashed boxes represent the coarse cells. Fine SFV cells (such as the two adjacent to #14 and #16) may be covered by a coarse PLG cell.
See Figure 4 for an illustration of Definition 5.1. When $\partial\Omega$ consists of multiple Jordan curves, we order PLG cells near each Jordan curve consecutively.
Lemma 5.2.
Suppose $\partial\Omega$ consists of only one single Jordan curve and the grid size $h$ is sufficiently small to resolve $\partial\Omega$. Then the total ordering in Definition 5.1 and the PLG discretization in Section 4.2 with $n=4$ yield
(32) $A_{22}=A_{22,c}+\begin{bmatrix}\mathbf{0}&A_{22,u}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}+\begin{bmatrix}\mathbf{0}&\mathbf{0}\\ A_{22,l}&\mathbf{0}\end{bmatrix}$,
where $\mathbf{0}$ represents a block matrix whose elements are all zero; the bandwidth of $A_{22,c}$ is at most 17, so are the dimensions of the square blocks $A_{22,u}$ and $A_{22,l}$.
Proof 5.3.
By $n=4$ and Definition 4.1, an SFV cell is at the center of a 5-by-5 box of regular cells. A nonempty cut-cell is either an SFV cell or a PLG cell and each row of $A_{22}$ corresponds to a PLG cell. Therefore, the distance from the center of a PLG cell to $\gamma$ is at most $\frac{5\sqrt{2}}{2}h$. In the worst case, the set $K$ of feasible nodes in Definition 4.2 is a 5-by-5 box, of which the starting point $\mathbf{q}$ is at the box corner. By the total ordering in Definition 5.1, the difference between the numbering of $\mathbf{q}$ and that of any multi-index in the box is bounded by $5\times\frac{5\sqrt{2}}{2}\approx 17.7$.
The above argument does not hold in a local neighborhood of $\gamma(0)$, where the difference of the numbering of two such PLG cells might be close to the total number of PLG cells. However, the number of these pairs of PLG cells is $O(1)$ and these large differences in PLG cell numbering can be assimilated either into $A_{22,l}$ or into $A_{22,u}$, whose dimensions, by similar arguments as above, are at most 17.
Definition 5.4.
The LU factorization of $A_{22}$ in (30) is given by
(33) $A_{22}:=\begin{bmatrix}B&P\\ Q&S\end{bmatrix}=\begin{bmatrix}L_{B}&\mathbf{0}\\ Y&L_{S}\end{bmatrix}\begin{bmatrix}U_{B}&X\\ \mathbf{0}&U_{S}\end{bmatrix}=:L_{22}U_{22},$
[…]
  1. (a)
    Perform an LU factorization on $B\in\mathbb{R}^{m\times m}$ to get $B=L_{B}U_{B}$;
  2. (b)
    Solve $L_{B}X=P$ for $X\in\mathbb{R}^{m\times k}$ by $k$ forward substitutions;
  3. (c)
    Solve $YU_{B}=Q$ for $Y\in\mathbb{R}^{k\times m}$ by $k$ backward substitutions;
  4. (d)
    Perform an LU factorization on $S^{\prime}=S-YX\in\mathbb{R}^{k\times k}$ to get $S^{\prime}=L_{S}U_{S}$.
To examine the complexity of the above LU factorization, we need
Lemma 5.5.
Suppose $A\in\mathbb{R}^{m\times m}$ has an LU factorization $A=LU$ and the bandwidth of $A$ is $p$. Then the bandwidths of $L$ and $U$ are both $p$. In addition, the complexity of this factorization via Gaussian elimination is $O(mp^{2})$.
Proof 5.6.
The first conclusion follows from [12, Theorem 4.3.1]. In the $k$th step of the Gaussian elimination, all non-zero elements from the $(k+1)$th row to the $\min(k+p,m)$th row need to be annihilated. Therefore, the total number of floating-point operations in this LU factorization is $\sum_{k=1}^{m-1}2\min(p,m-k)\cdot\min(p+1,m-k+1)$, yielding a complexity of $O(mp^{2})$.
Theorem 5.7.
For $A_{22}\in\mathbb{R}^{m\times m}$ in (30), we have $m=\operatorname{O}(h^{-1})$ and the complexity of the LU factorization of $A_{22}$ in Definition 5.4 is also $\operatorname{O}(h^{-1})$.
Proof 5.8.
$m=\operatorname{O}(h^{-1})$ follows from $\partial\Omega$ being a set of codimension one in $\Omega_{R}$.
For steps in Definition 5.4, Lemmas 5.2 and 5.5 imply that the complexity of (a) is $\operatorname{O}(m)$, Lemma 5.2 dictates that the complexity of each of (b) and (c) is $\operatorname{O}(m)$ and that the dimension of $S$ is $O(1)$, and thus the complexity of (d) is also $\operatorname{O}(1)$.
5.2 A block smoother
A fixed point iteration for solving a linear system $A\mathbf{u}=\mathbf{b}$ is an iteration of the form $\mathbf{u}^{(k+1)}=T\mathbf{u}^{(k)}+\mathbf{c}$ where $\mathbf{u}^{(k)}$ is the $k$th iterate that approximates $\mathbf{u}$ while $T$ and $\mathbf{c}$ are functions of $A$ and $\mathbf{b}$ satisfying $\mathbf{u}=T\mathbf{u}+\mathbf{c}$.
The Jacobi iteration is a fixed point iteration with $T_{J}=I-D^{-1}A$, $\mathbf{c}_{J}=D^{-1}\mathbf{b}$, where $D$ is the diagonal part of $A$. The weighted Jacobi iteration is another fixed point iteration of the form
(34) $\mathbf{u}^{(k+1)}:=(1-\omega)\mathbf{u}^{(k)}+\omega\mathbf{u}_{*}=(I-\omega D^{-1}A)\mathbf{u}^{(k)}+\omega\mathbf{c}_{J},$
where $\mathbf{u}_{*}=T_{J}\mathbf{u}^{(k)}+\mathbf{c}_{J}$. Due to the indefiniteness of $A_{22}$, a direct application to (30) would result in divergence.
Exploiting the block structure of (30) and the optimal complexity of the LU factorization in Theorem 5.7, we propose
Definition 5.9.
The block smoother for the linear system (30) is a fixed point iteration of the form
(35) $\begin{bmatrix}\mathbf{u}_{1}\\ \mathbf{u}_{2}\end{bmatrix}^{(k+1)}=T_{\omega}\begin{bmatrix}\mathbf{u}_{1}\\ \mathbf{u}_{2}\end{bmatrix}^{(k)}+\begin{bmatrix}\omega D_{11}^{-1}&\mathbf{0}\\ -\omega U_{22}^{-1}L_{22}^{-1}A_{21}D_{11}^{-1}&U_{22}^{-1}L_{22}^{-1}\end{bmatrix}\begin{bmatrix}\mathbf{b}_{1}\\ \mathbf{b}_{2}\end{bmatrix}$,
where $L_{22}U_{22}=A_{22}$ is the LU factorization in (33), $D_{11}$ is the diagonal of $A_{11}$, and
$T_{\omega}:=\begin{bmatrix}I&\mathbf{0}\\ -U_{22}^{-1}L_{22}^{-1}A_{21}&\mathbf{0}\end{bmatrix}\begin{bmatrix}I-\omega D_{11}^{-1}A_{11}&-\omega D_{11}^{-1}A_{12}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}$.
To derive (35), we first apply the weighted Jacobi to the first equation in (30),
(36) $\mathbf{u}^{(k+1)}_{1}=(I-\omega D_{11}^{-1}A_{11})\mathbf{u}^{(k)}_{1}+\omega D_{11}^{-1}\left(\mathbf{b}_{1}-A_{12}\mathbf{u}_{2}^{(k)}\right),$
and then exploit the LU factorization in (33) to solve for $\mathbf{u}_{2}^{(k+1)}$, i.e.,
(37) $L_{22}U_{22}\mathbf{u}_{2}^{(k+1)}=\mathbf{b}_{2}-A_{21}\mathbf{u}^{(k+1)}_{1}$.
After one iteration of (35), the residue vector on PLG cells, according to (31), is $\mathbf{r}_{2}^{(k+1)}:=\mathbf{b}_{2}-A_{21}\mathbf{u}_{1}^{(k+1)}-A_{22}\mathbf{u}_{2}^{(k+1)}$. Then (37) implies $\mathbf{r}_{2}^{(k+1)}=\mathbf{0}$. In other words, we always have $\mathbf{r}_{2}^{(k+1)}=\mathbf{0}$ for any $\mathbf{b}$; this is the key design of Definition 5.9.
The block smoother will also be applied in Algorithms 2 and 3 to the residual equation $A\mathbf{e}=\mathbf{r}$. Then the residual vector $\mathbf{r}$ must be updated after each iteration.
In classical multigrid theory, the value of $\omega$ in (34) is determined by minimizing the supremum of the set of all damping factors for high-frequency modes. For the diagonally dominant matrix resulting from the second-order FD discretization of the Laplacian operator, it is known [4, p. 21] [34, p. 31] that the optimal value of $\omega$ for the weighted Jacobi is $\omega=\frac{2}{3}$ and $\omega=\frac{4}{5}$ on $(0,1)$ and $(0,1)^{2}$, respectively. However, setting $\omega=\frac{4}{5}$ in (35) leads to numerical divergence in our fourth-order FV discretization, even on the regular domain $(0,1)^{2}$. Hence the smoothing property of the block smoother in Definition 5.9 is affected not only by the irregular domain but also by the fourth-order FV discretization. As such, it is difficult to analytically derive the optimal value of $\omega$ in (35).
In this work, we set $\omega=0.5$, which, according to extensive numerical experiments, preserves the smoothing property of the block smoother in (35) and minimizes the spectral radius of the two-grid correction operator in (41), cf. Table 1.
5.3 A cut-cell V-cycle
Define a hierarchy of levels of cut cells
(38) $\mathfrak{C}_{\Omega}(h_{f},n_{l},\epsilon):=\left\{\mathsf{C}_{\epsilon}^{h}(\Omega):h=h_{f},\ldots,2^{n_{l}-1}h_{f}\right\},$
where $h_{f}$ is the size of the finest grid, $n_{l}$ the number of levels of multigrid, and each level $\mathsf{C}_{\epsilon}^{h}(\Omega)$ the output of Algorithm 1 with $(\Omega,\Omega_{R},h,\epsilon)$ as the input. By Section 4, we have, on each level, a linear system $A^{h}\mathbf{u}^{h}=\mathbf{b}^{h}$ in the block form of (30).
[…]
Input: $(A^{h},\mathbf{b}^{h})$: the linear system resulting from discretizing (1) on $\mathsf{C}_{\epsilon}^{h}$;
$\mathbf{u}^{h}$: the initial guess of $(A^{h})^{-1}\mathbf{b}^{h}$;
$(\nu_{1},\nu_{2})$: the smoothing parameters.
Side-effect: $\mathbf{u}^{h}$ is updated as a better approximation to $(A^{h})^{-1}\mathbf{b}^{h}$.
1:if $h$ is the grid size of the coarsest level then
2:  $\mathbf{u}^{h}\leftarrow\text{{BottomSolver}}(A^{h},\mathbf{b}^{h})$
3:else
4:  for $i=1,\ldots,\nu_{1}$ do
5:   $\mathbf{u}^{h}\leftarrow\text{{Smooth}}(A^{h},\mathbf{u}^{h},\mathbf{b}^{h})$ // see (35)
6:  end for
[…]
8:  $\mathbf{e}^{2h}\leftarrow\text{{V-cycle}}(A^{2h},\mathbf{0}^{2h},\mathbf{r}^{2h},\nu_{1},\nu_{2})$ // $\parbox[t]{7.11317pt}{}//\mbox{ the initial guess is a zero vector}$
[…]
10:  for $i=1,\ldots,\nu_{2}$ do
11:   $\mathbf{u}^{h}\leftarrow\text{{Smooth}}(A^{h},\mathbf{u}^{h},\mathbf{b}^{h})$ // see (35)
12:  end for
13:end if
We present in Algorithm 2 a cut-cell V-cycle that appears very similar to standard geometric multigrid V-cycles. At line 2, we directly solve the linear system if the current grid is the coarsest one. Otherwise, we use (35) to block-smooth $\mathbf{u}^{h}$ $\nu_{1}$ times at lines 4–6, restrict the corresponding residual to the next coarser level at line 7, call Algorithm 2 recursively to solve the residual equation on the coarser level at line 8, correct the solution by the error interpolated from the coarse level at line 9, and finally block-smooth $\mathbf{u}^{h}$ $\nu_{2}$ times at lines 10–12.
For the restriction operator at line 7, we first observe that each irregular cell is a PLG cell and hence its residual becomes zero after one block smoothing. Furthermore, if an irregular fine cell is covered by some coarse cell, then all fine cells (regular or irregular) covered by this coarse cell have their residuals as identically zero after one round of block smoothing, due to the fact of the refinement ratio being 2 and the width of SFV stencil being 5; see Figure 4. Consequently, residual restriction only happens between regular fine cells and regular coarse cells. These observations obviate the need of volume weighting in residual restriction and lead to a restriction operator $I_{h}^{2h}:\mathbf{r}^{h}\rightarrow\mathbf{r}^{2h}$ of the simple form
(39) $\left\langle r^{2h}\right\rangle_{\lfloor\frac{\mathbf{i}}{2}\rfloor}=2^{-{\scriptsize\textsf{D}}}\sum\nolimits_{\mathbf{j}\in\left\{0,1\right\}^{\scriptsize\textsf{D}}}\left\langle r^{h}\right\rangle_{\mathbf{i}+\mathbf{j}},$
where $\lfloor\mathbf{k}\rfloor$ is the greatest multi-index less than or equal to the D-tuple $\mathbf{k}$ of real numbers. Thanks to the FV formulation, (39) incurs no discretization errors. On the other hand, the interpolation operator $I_{2h}^{h}:\mathbf{e}^{2h}\rightarrow\mathbf{e}^{h}$ is given by
(40) $\left\langle e^{h}\right\rangle_{\mathbf{i}}=\left\langle e^{2h}\right\rangle_{\lfloor\frac{\mathbf{i}}{2}\rfloor}$
so that (39) and (40) satisfy the variational property $I_{2h}^{h}=2^{{\scriptsize\textsf{D}}}(I_{h}^{2h})^{\top}$.
Residual restriction to a coarse regular cell might involve both PLG fine cells and SFV fine cells; for example, the two PLG fine cells numbered #14 and #16 in Figure 4 and the two SFV cells to the right of them are covered by a PLG coarse cell, whose residual vanishes after a single pre-smoothing. Similarly, after errors on coarse PLG cells are interpolated to fine cells, those of fine PLG cells are immediately annihilated by one around of post-smoothing. In addition, each fine PLG cell is covered by some coarse PLG cell. These observations, together with the classical theory of geometric multigrid, furnish strong heuristics in supporting the convergence of the cut-cell V-cycle. They also suggest that both $\nu_{1}$ and $\nu_{2}$ be at least 1.
For the particular case of $n_{l}=2$ in (38), the cut-cell V-cycle reduces to a two-grid correction operator [4, p. 82] given by
(41) $TG:=T_{\omega}^{\nu_{2}}\left[I-I_{2h}^{h}(A^{2h})^{-1}I_{h}^{2h}A^{h}\right]T_{\omega}^{\nu_{1}}.$
Table 1: Values of $\rho(TG)$, the spectral radius of $TG$ in (41) with $\omega=\frac{1}{2}$, for elliptic problems on various domains as specified in Section 6. In particular, the elliptic equation solved on the rotated square in Figure 6(b) has a cross-derivative term and the irregular boundary in Figure 8(a) is equipped with a Neumann boundary condition. For each case, we select three successively refined grids so that the most significant digits of the calculated spectral radii are the same on the two finest grids. The pairs of integers in the first row are values of $(\nu_{1},\nu_{2})$.
Test cases $(1,0)$ $(1,1)$ $(2,1)$ $(2,2)$ $(3,3)$
the unit square $(0,1)^{2}$ in Figure 6(a) 1.080 0.758 0.603 0.513 0.378
the rotated square $\Omega_{r}$ in Figure 6(b) 1.110 0.745 0.523 0.421 0.275
$(0,1)^{2}$ minus a flower in Figure 7(a) 1.069 0.698 0.483 0.414 0.283
$(0,1)^{2}$ minus four disks in Figure 8(a) 1.272 0.878 0.641 0.488 0.308
We numerically calculate the spectral radii $\rho(TG)$ of $TG$ in (41), also known as the convergence factor of $TG$, for the test problems in Section 6, verify the independence of $\rho(TG)$ on $h$ for each test case, and collect their values in Table 1. Before these results are discussed, we mention the result in [34, Section 4.6.1] that 0.084 is the value of the convergence factor of the classical two-grid operator with $(\nu_{1},\nu_{2})=(2,2)$, Gauss-Seidel smoothing, full weighting restriction, and bilinear interpolation for second-order FD discretization of Poisson’s equation in the unit square.
[…]
For $\nu_{1}=\nu_{2}=2$, all values of $\rho(TG)$ are less than 0.52. Then it follows from $0.52^{3.79}\approx 0.084$ that, to obtain the same ratio of residual reduction, the number of cut-cell V-cycles needs to be 3.79 times as many as that of classical multigrid V-cycles. Fortunately, this gap can be very much reduced by bringing a cut-cell FMG cycle into the big picture.
5.4 A cut-cell FMG cycle
The convergence factor $\rho$ of a V-cycle is usually independent of the grid size $h$ and is less than 1, and thus it takes $\operatorname{O}(\log(h^{-1}))$ V-cycles to solve the linear system $A^{h}\mathbf{u}^{h}=\mathbf{b}^{h}$. It is well known from the multigrid literature [4, p. 77–78] that this suboptimal complexity of $\operatorname{O}(\log(h^{-1}))$ V-cycles can be improved to the optimal complexity of $\operatorname{O}(1)$ FMG cycles.
[…]
Input: $(A^{h},\mathbf{r}^{h})$: a residual equation corresponding to the linear system (29);
$(\nu_{1},\nu_{2})$: the smoothing parameters.
Output: An approximation to $(A^{h})^{-1}\mathbf{r}^{h}$.
1:if $h$ is the grid size of the coarsest level then
2:  return $\text{{BottomSolver}}(A^{h},\mathbf{r}^{h})$
3:end if
[…]
8:return $\mathbf{e}^{h}$
[…]
[width=0.6]tikz/FMGCycle
Figure 5: Illustrating the FMG cycle in Algorithm 3 on a hierarchy of four levels. FMG begins with a descent at line 4 to the coarsest level $\mathsf{C}_{\epsilon}^{8h}$; this is represented by the first three downward dashed line segments. Then the solution is interpolated to $\mathsf{C}_{\epsilon}^{4h}$ at line 6 and used as the initial guess to the V-cycle at line 7 on $\mathsf{C}_{\epsilon}^{4h}$. This “interpolation ++ V-cycle” process is repeated recursively: the interpolation is represented by an upward dashed line and the V-cycles are represented by the solid lines. This FMG cycle is more effective than the V-cycle because it comes up with a much better initial guess, cf. line 8 at Algorithm 2.
Finally, we claim that the cut-cell FMG cycle in Algorithm 3 is of the optimal complexity $\operatorname{O}(h^{-2})$. In setting up the block smoother, $A_{11}$ is initialized in $O(h^{-2})$ time while all other block matrices are computed in $O(h^{-1})$ time, cf. Theorem 5.7. In solving $A^{h_{f}}\mathbf{u}^{h_{f}}=\mathbf{b}^{h_{f}}$ on $\mathfrak{C}_{\Omega}(h_{f})$, the entire computation cost of an FMG cycle is $O(h^{-2})$, the same as that of the deepest V-cycle, because an FMG cycle in 2D is at most $\frac{4}{3}$ times more expensive than the deepest V-cycle [4, p. 47–48].
6 Numerical tests
In this section we demonstrate the fourth-order accuracy and the optimal efficiency of our cut-cell geometric multigrid method by results of various test problems. To facilitate accuracy comparisons of our method to the second- and fourth-order EB methods in [17, 9], we follow [17, 9] to measure computational errors by the $L^{p}$ norms,
(42) $\|u\|_{p}=\begin{cases}\left(\frac{1}{\|\Omega\|}\sum_{{\cal C}_{\mathbf{i}}\in\mathsf{C}_{\epsilon}^{h}(\Omega)}\left\|\mathcal{C}_{\mathbf{i}}\right\|\cdot\left|\left\langle u\right\rangle_{\mathbf{i}}\right|^{p}\right)^{\frac{1}{p}}&\text{if }p=1,2;\\ \max_{{\cal C}_{\mathbf{i}}\in\mathsf{C}_{\epsilon}^{h}(\Omega)}\left|\left\langle u\right\rangle_{\mathbf{i}}\right|&\text{if }p=\infty,\end{cases}$
where $\mathsf{C}_{\epsilon}^{h}(\Omega)$ is the set of nonempty cut cells in (9).
6.1 A rotated square
(a) Results on the unit square
(b) Results on the rotated unit square $\Omega_{r}$
Figure 6: Results of the proposed cut-cell multigrid method in solving the two equivalent tests in Section 6.1 with $h=\frac{1}{256}$. The two exact solutions are related by a rotation of $\frac{\pi}{6}$ around $(0,0)$.
Table 2: Error norms and convergence rates of the proposed method with $\epsilon=0.1$ for solving the tests in Section 6.1. $\Omega_{r}$ is obtained by rotating $(0,1)^{2}$ around the origin by $\frac{\pi}{6}$; see Figure 6(b).
$\Omega$ $h=\frac{1}{64}$ rate $h=\frac{1}{128}$ rate $h=\frac{1}{256}$ rate $h=\frac{1}{512}$
$(0,1)^{2}$ $L^{\infty}$ 3.68e-08 4.00 2.30e-09 4.00 1.44e-10 3.99 9.02e-12
$L^{1}$ 1.13e-08 4.01 7.00e-10 4.01 4.35e-11 3.91 2.89e-12
$L^{2}$ 1.50e-08 4.01 9.32e-10 4.00 5.81e-11 3.97 3.71e-12
$\Omega_{r}$ $L^{\infty}$ 1.35e-07 3.93 8.85e-09 3.93 5.79e-10 3.95 3.75e-11
$L^{1}$ 4.83e-08 4.03 2.95e-09 3.91 1.96e-10 3.91 1.31e-11
$L^{2}$ 5.92e-08 3.99 3.72e-09 3.89 2.51e-10 3.91 1.67e-11
This test consists of two cases. First, we set $\Omega=(0,1)^{2}$ and $(a,b,c)=(1,0,2)$ in (1), for which the exact solution is
(43) $\forall(x_{1},x_{2})\in\overline{\Omega},\quad u(x_{1},x_{2})=\sin(4x_{1})\cos(3x_{2}),$
and the boundary condition is the Dirichlet condition from (43). Due to the regularity of $\Omega$, all cut cells in $\mathsf{C}_{\epsilon}^{h}(\Omega)$ are SFV cells, the blocks $A_{21}$, $A_{12}$, and $A_{22}$ vanish, and the linear system (30) reduces to that of the standard fourth-order FV discretization of (1). Also, the block smoother in Definition 5.9 reduces to the weighted Jacobi.
In the second case, the domain $\Omega_{r}$ is obtained by rotating the unit square around the origin by $\frac{\pi}{6}$; see Figure 6(b). For $\left(a,b,c\right)=\frac{1}{4}\left(5,-2\sqrt{3},7\right)$, the exact solution $u:\overline{\Omega}\to\mathbb{R}$ of (1) is obtained by rotating that in (43) by $\frac{\pi}{6}$. A Dirichlet condition is imposed to ensure that the only difference of these two cases is the regularity of the boundary. The main goal of this setup is to examine how the cross-derivative term and the PLG discretizations affect the solution errors.
For the two equivalent systems, numerical solutions with $h=\frac{1}{256}$ are shown in Figure 6(b), with error norms and convergence rates listed in Table 2, where the fourth-order accuracy are clearly demonstrated. Each error norm on the irregular domain is greater than its counterpart on the regular domain, due to the PLG discretization and the larger SFV stencil for the additional cross-derivative term. However, the ratio of the two error norms is bounded by 4.6 and we consider the slightly lower accuracy as a reasonable cost for PLG and the cross-derivative term.
6.2 A square minus a flower
(a) the numerical solution
(b) the truncation error
(c) the solution error
Figure 7: Results of the proposed method in solving the problem in Section 6.2 with $h=\frac{1}{80}$. Different from subplots (a,c), subplot (b) has a logarithmic scale for representing truncation errors.
Table 3: Truncation and solution errors of the proposed cut-cell method with $\epsilon=0.02$ and a second-order EB method [17] for solving the test problem in Section 6.2.
Truncation errors of the EB method by Johansen and Colella [17]
$h=\frac{1}{40}$ rate $h=\frac{1}{80}$ rate $h=\frac{1}{160}$ rate $h=\frac{1}{320}$
$L^{\infty}$ 1.66e-03 2.0 4.15e-04 2.0 1.04e-04 2.0 2.59e-05
Truncation errors of the proposed fourth-order cut-cell method
$h=\frac{1}{40}$ rate $h=\frac{1}{80}$ rate $h=\frac{1}{160}$ rate $h=\frac{1}{320}$
$L^{\infty}$ 6.53e-04 3.02 8.03e-05 2.49 1.42e-05 3.28 1.47e-06
$L^{1}$ 1.48e-05 4.01 9.23e-07 4.02 5.68e-08 4.11 3.29e-09
$L^{2}$ 5.27e-05 3.60 4.35e-06 3.52 3.79e-07 3.64 3.05e-08
Solution errors of the EB method by Johansen and Colella [17]
$h=\frac{1}{40}$ rate $h=\frac{1}{80}$ rate $h=\frac{1}{160}$ rate $h=\frac{1}{320}$
$L^{\infty}$ 4.78e-05 1.85 1.33e-05 1.98 3.37e-06 1.95 8.72e-07
Solution errors of the proposed fourth-order cut-cell method
$h=\frac{1}{40}$ rate $h=\frac{1}{80}$ rate $h=\frac{1}{160}$ rate $h=\frac{1}{320}$
$L^{\infty}$ 5.42e-07 4.98 1.72e-08 3.73 1.29e-09 3.68 1.01e-10
$L^{1}$ 7.72e-08 5.23 2.06e-09 3.86 1.42e-10 3.84 9.90e-12
$L^{2}$ 1.31e-07 5.20 3.54e-09 3.89 2.39e-10 3.84 1.66e-11
In this test, we follow [17, Problem 3] to solve Poisson’s equation on an irregular domain $\Omega=R\cap\Omega_{1}$, where $R=(-0.5,0.5)^{2}$, $\Omega_{1}=\left\{\left(r,\theta\right):r>0.25+0.05\cos 6\theta\right\}$, and $\left(r,\theta\right)$ are the polar coordinates satisfying $(x_{1},x_{2})=(r\cos\theta,r\sin\theta)$. As shown in Figure 7(a), we set the exact solution as
(44) $\forall(x_{1},x_{2})\in\overline{\Omega},\quad u(x_{1},x_{2})=u(r,\theta)=r^{4}\cos 3\theta$
and impose Dirichlet and Neumann conditions on $\partial R$ and $\partial\Omega_{1}$, respectively.
Due to the symmetric FV formulas in Section 4.1, the truncation error $\tau_{\mathbf{i}}$ for an SFV cell ${\cal C}_{\mathbf{i}}$ is $O(h^{4})$. For a PLG cell ${\cal C}_{\mathbf{i}}$, however, it follows from (21) and the opening paragraph of Section 4.3 that the truncation error for the $\mathbf{i}$th cut cell is given by $\tau_{\mathbf{i}}:={\beta}_{\min}^{\top}\overline{\mathbf{u}}-\left\langle\mathcal{L}u\right\rangle_{\mathbf{i}}=O(h^{3})$. This is confirmed both in Figure 7(b) and Table 3, where the convergence rates of truncation errors are asymptotically close to 3, 3.5, and 4 in the $L^{\infty}$, $L^{2}$, and $L^{1}$ norms, respectively. In Figure 7(c), the non-uniformness of truncation errors causes solution errors to be oscillatory; however, the magnitude of solution errors is very small ($\sim 10^{-10}$) even for the large grid size $h=\frac{1}{80}$. More importantly, the large truncation errors near the boundary do not affect the fourth-order accuracy of solution errors; this is well known for FD/FV methods and is confirmed in Table 3.
Truncation errors and solution errors of the classical second-order EB method by Johansen and Colella [17] are also listed in Table 3. Clearly, our method is much more accurate: the $L^{\infty}$ solution error of our method on the coarest grid of $h=\frac{1}{40}$ is smaller than that of the second-order EB method on the finest grid of $h=\frac{1}{320}$.
6.3 A square minus four disks
Consider a problem in [9, §5.2] of solving Poisson’s equation on the domain $\Omega=R\setminus\Omega_{d}$, where $R=(0,1)^{2}$ and $\Omega_{d}$ is the closure of the union of four disks, whose centers and radii $(c_{1},c_{2};r)$ are $(0.5,0.5;0.2)$, $(0.5,0.735;0.1)$, $(0.2965,0.3825;0.1)$, and $(0.7035,0.3825,0.1)$. At each of the six kinks on $\Omega_{d}$, a level-set function that implicitly represents $\partial\Omega_{d}$ would be $C^{1}$ discontinuous. Following [9, §5], we set the exact solution as
(45) $\forall(x_{1},x_{2})\in\overline{\Omega},\quad u=\sin(\pi x_{1})\sin(\pi x_{2})$
and impose on $\partial R$ a Dirichlet condition from (45).
In Table 4, error norms and convergence rates of our cut-cell method and the fourth-order EB method in [9] are presented for solving this test with Dirichlet and Neumann conditions on $\partial\Omega_{d}$. The fourth-order EB method performs poorly: its converge rates barely reach 2 and 1 for Dirichlet and Neumann conditions, respectively; as shown in [9, Fig. 8], its largest solution errors concentrate around the six kinks. This is not surprising because, as discussed in Section 1, the error of the normal vector near a kink is $O(1)$ and thus the integral of fluxes over faces of an irregular cut cell is calculated with an error of $O(h)$. At the presence of kinks, this accuracy deterioration is unavoidable if the discretization of (1) is coupled with an implicit representation of the domain boundary. Although the fourth-order accuracy can be recovered by smoothing the geometric description, this mollification process requires substantial extra care and its effectiveness depends largely on mollification formulas and the nature of the governing equation [9, §6].
Table 4: Solution errors and convergence rates of the proposed cut-cell method with $\epsilon=0.08$ and a fourth-order EB method [9] in solving the test problem in Section 6.3.
$h=\frac{1}{64}$ rate $h=\frac{1}{128}$ rate $h=\frac{1}{256}$ rate $h=\frac{1}{512}$
4th-order EB method without mollifying kinks; a Dirichlet condition on $\partial\Omega_{d}$
$L^{\infty}$ 2.80e-03 3.05-3.05 2.32e-02 3.10 2.71e-03 2.06 6.50e-04
$L^{1}$ 3.23e-05 2.82 4.56e-06 4.49 2.01e-07 1.76 5.94e-08
$L^{2}$ 1.09e-04 0.56 7.40e-05 3.90 4.97e-06 1.13 2.26e-06
4th-order EB method with kink mollification; a Dirichlet condition on $\partial\Omega_{d}$
$L^{\infty}$ 2.64e-08 4.06 1.58e-09 4.03 9.65e-11 3.85 6.67e-12
$L^{1}$ 1.08e-08 4.08 6.38e-10 4.03 3.88e-11 3.84 2.69e-12
$L^{2}$ 1.34e-08 4.11 7.76e-10 4.05 4.68e-11 3.86 3.23e-12
the proposed cut-cell method; a Dirichlet condition on $\partial\Omega_{d}$
$L^{\infty}$ 5.44e-08 5.48 1.22e-09 3.88 8.28e-11 4.52 3.62e-12
$L^{1}$ 9.50e-09 4.87 3.26e-10 4.41 1.54e-11 4.19 8.41e-13
$L^{2}$ 1.17e-08 4.94 3.81e-10 4.41 1.79e-11 4.19 9.84e-13
$h=\frac{1}{64}$ rate $h=\frac{1}{128}$ rate $h=\frac{1}{256}$ rate $h=\frac{1}{512}$
4th-order EB method without mollifying kinks; a Neumann condition on $\partial\Omega_{d}$
$L^{\infty}$ 3.10e-02 0.14 2.82e-02 0.52 1.97e-02 0.38 1.52e-02
$L^{1}$ 1.42e-03 0.58-0.58 2.13e-03 0.62 1.38e-03 1.03 6.75e-04
$L^{2}$ 2.72e-03 0.24-0.24 3.20e-03 0.67 2.02e-03 0.96 1.04e-03
4th-order EB method with kink mollification; a Neumann condition on $\partial\Omega_{d}$
$L^{\infty}$ 1.84e-07 4.02 1.13e-08 3.97 7.21e-10 3.74 5.39e-11
$L^{1}$ 5.83e-08 3.94 3.78e-09 3.97 2.40e-10 3.86 1.65e-11
$L^{2}$ 7.35e-08 3.96 4.73e-09 3.98 2.99e-10 3.87 2.04e-11
the proposed 4th-order cut-cell method; a Neumann condition on $\partial\Omega_{d}$
$L^{\infty}$ 2.07e-07 4.29 1.06e-08 4.01 6.56e-10 3.79 4.76e-11
$L^{1}$ 2.80e-08 5.02 8.65e-10 3.95 5.58e-11 3.79 4.04e-12
$L^{2}$ 3.98e-08 4.63 1.61e-09 4.09 9.43e-11 3.75 7.03e-12
(a) A Neumann condition on the disks $\partial\Omega_{d}$
(b) A Dirichlet condition on the disks $\partial\Omega_{d}$
Figure 8: Solution errors of the proposed cut-cell method in solving the problem in Section 6.3 with $h=\frac{1}{512}$. A Dirichlet condition is applied on the square boundary. All boundary conditions are derived from (45).
In comparison, convergence rates of our cut-cell method are closed to 4 in both cases and its solution errors in all norms are smaller than those of the fourth-order EB method with mollifications. As shown in Figure 8, solution errors of our cut-cell method are not concentrated at the six kinks. This is also unsurprising because (i) the explicit representation of domain boundary by cubic splines admits a fourth- and higher-order approximation of the geometry of any irregular cut cell and (ii) the integrals of solutions over an irregular cut cell can be approximated to very high-order accuracy by Green’s theorem and Gauss quadrature formulas. In summary, the integral formulation of our cut-cell method (enabled by explicit representation of geometry) is advantageous over the differential formulation of previous EB methods.
6.4 A panda
To showcase the capability of the proposed cut-cell method in handling complex topology and geometry, we numerically solve (1) on the domain of a panda shown in Figure 9(a), which is adapted from that in [40, Figure 10] with a sufficent number of breakpoints. The same spline representation of the panda boundary is used for all grid sizes. The complex topology and geometry of the panda pose significant challenges to a numerical solver.
(a) the numerical solution
(b) the solution error
Figure 9: Results of the proposed cut-cell method in solving the panda problem in Section 6.4 with $(a,b,c)=(1,1,2)$ and $h=\frac{1}{512}$.
Table 5: Solution errors and convergence rates of the proposed cut-cell method with $\epsilon=0.01$ in solving the panda problem in Section 6.4, where we impose the Dirichlet condition (46) on the boundary.
$(a,b,c)$ $h=\frac{1}{256}$ rate $h=\frac{1}{512}$ rate $h=\frac{1}{1024}$ rate $h=\frac{1}{2048}$
$(1,1,2)$ $L^{\infty}$ 7.97e-05 5.05 2.41e-06 4.03 1.48e-07 4.02 9.14e-09
$L^{1}$ 5.00e-06 3.97 3.18e-07 3.96 2.05e-08 3.97 1.31e-09
$L^{2}$ 7.55e-06 3.96 4.87e-07 3.96 3.13e-08 3.98 1.99e-09
$(1,0,2)$ $L^{\infty}$ 5.47e-05 4.58 2.29e-06 4.01 1.42e-07 4.01 8.82e-09
$L^{1}$ 4.75e-06 3.94 3.09e-07 3.95 2.00e-08 3.97 1.28e-09
$L^{2}$ 7.25e-06 3.93 4.76e-07 3.95 3.08e-08 3.97 1.96e-09
The exact solution of this test is
(46) $\forall(x_{1},x_{2})\in\overline{\Omega},\quad u(x_{1},x_{2})=\left(x_{1}^{2}+x_{2}^{2}-1\right)[\sin(50x_{1})+\cos(50x_{2})].$
We impose the corresponding Dirichlet condition on the boundary of the panda and numerically solve (1) for $(a,b,c)=(1,1,2)$ and $(1,0,2)$. Results of our cut-cell method for the case with the cross-derivative term on the grid of $h=\frac{1}{512}$ are plotted in Figure 9 and the error norms are listed in Table 5, where the convergence rates are very close to 4.0 in all norms, demonstrating the fourth-order accuracy and the capability of our method in handling complex domains. In addition, quantitative results for $(a,b,c)=(1,1,2)$ are very close to those for $(a,b,c)=(1,0,2)$, indicating that the cross-derivative term is handled satisfactorily.
Table 6: CPU time in seconds for solving the panda test in Section 6.4 with $(a,b,c)=(1,1,2)$ on an AMD Threadripper PRO 3975WX at 4.0Ghz. For each of the four setup steps, we also report its percentage of the entire cost of setup in a pair of parentheses.
Stages steps $h=\frac{1}{512}$ $h=\frac{1}{1024}$ $h=\frac{1}{2048}$
Setup generate the set $\mathsf{C}_{\epsilon}^{h}(\Omega)$ of cut cells by Algorithm 1 0.081 (3.7%) 0.303 (6.7%) 1.12 (11.7%)
locate a poised lattice for each PLG cell 0.087 (3.9%) 0.180 (4.0%) 0.407 (4.2%)
determine the linear system in (30) by steps in Section 4 2.01 (90.8%) 3.99 (87.8%) 7.94 (82.3%)
compute the block smoother in Definition 5.9 0.036 (1.6%) 0.073 (1.5%) 0.176 (1.8%)
the entire cost of setup, i.e., the sum of the above four steps 2.21 (100%) 4.55 (100%) 9.64 (100%)
Solve block smoothing only 0.521 2.38 9.84
the entire cost of FMG cycles 1.08 3.63 14.3
Computational costs of the main components of our cut-cell method are reported in Table 6 for the case of $(a,b,c)=(1,1,2)$. The generation of cut cells by Algorithm 1 clearly has the $O(h^{-2})$ complexity. In contrast, all other setup steps have the optimal $O(h^{-1})$ complexity, confirming the analysis in Section 4 and Sections 5.1 and 5.2. In particular, the complexity of determining the linear system (30) is only $O(h^{-1})$ because the block $A_{11}$ is never assembled but applied “on the fly” inside the weighted Jacobi. As indicated by the last row of Table 6, FMG cycles have the optimal complexity of $O(h^{-2})$, which confirms our analysis in Section 5.4.
The cost of generating cut cells is very much dominated by that of determining the linear system (30), which holds even on the finest grid. Consequently, the cost of the entire initial setup displays a roughly linear growth as the grid size $h$ is reduced. Being the most expensive component of V-cycles and FMG cycles, block smoothing consumes more CPU time than the initial setup on the finest grid, since the $O(h^{-2})$ growth of its cost is higher than the linear growth.
6.5 FMG efficiency
The panda is a good representative of complex domains while the rotated square in Figure 6(b) that of the other extreme of irregular but simple domains. For the rotated square in Figure 6(b) with $h=\frac{1}{256},\frac{1}{512},\frac{1}{1024}$, we record computational costs (not shown) of the main components. The cost of generating cut cells grows quadratically with respect to the reduction of $h$ while those of all other steps in the initial setup grows linearly; the $O(h^{-2})$ complexity of FMG cycles are also confirmed. In addition, the consumed CPU time in seconds for $h=\frac{1}{1024}$ is 1.89, 2.62, 3.57, and 5.31 for the determination of (30), the entire initial setup, the block smoothing, and all FMG cycles, respectively. The cost ratio of block smoothing over all FMG cycles is $\frac{3.57}{5.31}\approx 0.67$ for the rotated square, which is very close to that ($\frac{2.38}{3.63}\approx 0.66$) for the panda, cf. Table 6. Due to the simple geometry of the rotated square, the cost ratio of the entire setup over all FMG cycles, $\frac{2.62}{5.31}\approx 0.49$, is much smaller than that ($\frac{4.55}{3.63}\approx 1.25$) for the panda test. We sum up main conclusions of the above discussions as follows.
  • The proposed cut-cell method has the optimal complexity of $O(h^{-2})$.
  • On a coarse grid, the initial setup might be more expensive than the FMG cycles. However, there exists a grid size $h^{*}$ such that, for any $h<h^{*}$, the computational cost of the initial setup is less than that of the FMG cycles.
  • Block smoothing consumes $\frac{2}{3}$ of the entire CPU time of FMG cycles and is asymptotically the most expensive component of the proposed method.
(a) Section 6.1 with $h=\frac{1}{512}$
(b) Section 6.4 with $h=\frac{1}{2048}$
Figure 10: Performance of FMG cycles of our method with $\nu_{1}=\nu_{2}=3$ for tests on the unit square $(0,1)^{2}$, the rotated square $\Omega_{r}$, and the panda in Sections 6.1 and 6.4. The ordinates are the relative residuals/errors in $L^{\infty}$ norm, the abscissa is the iteration number of FMG cycles. The condition number of an FMG cycle is indicated by the smallest solution error that remains constant under more multigrid iterations.
Table 7: Averaged reduction rates of residuals in solving problems in Sections 6.1, 6.2, 6.3 and 6.4.
Tests cycles $(\nu_{1},\nu_{2})=(2,1)$ $(\nu_{1},\nu_{2})=(2,2)$ $(\nu_{1},\nu_{2})=(3,3)$
the unit square $(0,1)^{2}$ V 0.858 0.410 0.330
FMG 0.672 0.246 0.131
the rotated square $\Omega_{r}$ V 0.474 0.266 0.216
FMG 0.376 0.162 0.103
$(0,1)^{2}$ minus a flower V 0.384 0.318 0.221
FMG 0.296 0.207 0.108
$(0,1)^{2}$ minus four disks (a Dirichlet condition) V 0.298 0.236 0.161
FMG 0.292 0.120 0.063
$(0,1)^{2}$ minus four disks (a Neumann condition) V 0.373 0.336 0.279
FMG 0.347 0.208 0.135
panda with $(a,b,c)=(1,0,2)$ V 0.652 0.357 0.178
FMG 0.548 0.233 0.119
panda with $(a,b,c)=(1,1,2)$ V 0.661 0.359 0.189
FMG 0.636 0.270 0.126
In Figure 10, we show the performance of FMG cycles of our method in solving the tests in Sections 6.1 and 6.4. For the unit and rotated squares, each FMG cycle respectively reduces the residual by a factor of 7.6 and 9.7. As for the panda tests, each FMG cycle reduces the residual by a factor of 8.4 and 7.9, respectively. These reduction rates confirm the discussions on Table 1 in Section 5.3 and are comparable to those of classical geometric multigrid methods.
Finally in Table 7, we list averaged reduction rates of our multigrid cycles for problems in Sections 6.1, 6.2, 6.3 and 6.4. The improvement of FMG cycles over V-cycles are clearly demonstrated, verifying the claim in the ending sentence of Section 5.3. Altogether, Tables 6 and 7 imply that the choice of $(\nu_{1},\nu_{2})=(3,3)$ is more cost-effective than those of $(\nu_{1},\nu_{2})=(2,1)$ and $(\nu_{1},\nu_{2})=(2,2)$. For complex domains and moderate grid sizes, it might be appropriate to choose even greater values of $(\nu_{1},\nu_{2})$.
7 Conclusions
We have proposed a fourth-order cut-cell multigrid method for solving constant-coefficient elliptic equations on 2D irregular domains with the optimal complexity of $O(h^{-2})$. Based on the Yin space, our method is able to handle arbitrarily complex topology and geometry. Results of comprehensive numerical tests demonstrate the accuracy, efficiency, robustness, and generality of the new method.
Prospects for future research are as follows. First, this work motivates theoretical investigations on the effectiveness of the proposed multigrid method. Second, we will augment the proposed cut-cell method to elliptic equations with variable coefficients. Lastly, we will follow the GePUP formulations in [38, 21] to develop a fourth-order INSE solver on irregular domains, for which the proposed method in this work can be reused to solve pressure Poisson equations and Helmholtz-like equations.
Acknowledgments. We acknowledge helpful comments from Shaozhen Cao, Lei Pang, and Chenhao Ye, graduate students at the school of mathematical sciences in Zhejiang University.